# BigTokDetect: A Clinically-Informed Vision–Language Modeling Framework for Detecting Pro-Bigorexia Videos on TikTok

### v2

@@ B.1 Annotator Profiles
We enlist 16 subject-matter experts, ranging from clinical psychologists and psychiatrists to computational social scientists, for video annotation. Our annotators are research collaborators and co-authors who volunteered their clinical expertise with full knowledge of the project’s scope and scientific objectives, ensuring informed participation in this sensitive content annotation task. Table 11 details each annotator’s ID, batch assignment, area of expertise, and gender.
@@ B.3 Annotation Platform
The HTML interface (Figure 2) presents each TikTok video with its caption and hashtags (if any). The form beneath the video prompts the annotator to mark if the video is: (1) Showing One’s Body (yes/no); includes (2) Commercial Sponsorship (yes/no); (3) 2 boxes for Type of harm (1 mandatory, 1 optional), consisting of primary categories and subcategories, plus “Relevant but Not Listed,” “Irrelevant,” and “Unsure—Needs Consultation” options; and (4) Severity of Harm on a discrete scale ranging from 1 (Not Harmful) to 5 (Extremely Harmful). A text box allows annotators to write additional comments and edge-case observations.
@@ B.4 Annotation Processing
After Round 2, we apply a five-step pipeline to prepare the data for consensus review and final curation:
@@
    Data Filtering: remove videos that are unavailable (deleted) or non-English, as indicated in annotators’ notes.
@@
    Reliability Assessment: compute Cohen’s $\kappa$ for both strong (exact primary–secondary match) and weak (primary-only match) agreement to quantify inter-rater reliability.
@@
    Consensus Review and Re-annotation: filter out all items by agreement level (strong, weak, disagreement), sample a subset of weak and disagreement cases for each annotator pair, hold a batch-specific meeting to discuss and jointly re-annotate selected cases to reach consensus, then return the remaining weak and disagreement items to annotators for independent re-annotation.
@@ B.5 Annotation Statistics
Inter-annotator agreement substantially improves from Round 2 to Final, with Cohen’s $\kappa$ values increasing from moderate agreement (0.43-0.69 strong $\kappa$, 0.59-0.83 weak $\kappa$) to good-to-excellent agreement (0.58-0.81 strong $\kappa$, 0.78-0.94 weak $\kappa$). The consistently high weak $\kappa$ values (>0.9 in most batches) in the Final round indicate that annotators achieved near-perfect agreement on the broader annotation categories.
@@
Table 6: Cohen’s kappa coefficients for the inter-agreement rate among annotators in Rounds 2 and 3
@@
(b) Subcategories (reordered)
Table 7: Distribution of Annotated Primary Categories and Subcategories
@@ Appendix C Data
Our dataset comprises TikTok videos annotated for hierarchical classification across two levels of granularity. The primary category classification categorizes content into six broad categories related to body image and health behaviors, while the subcategory classification provides fine-grained classification into 16 specific subcategories. Severity score scale define the intensity of harm emotion. The severity scores are most heavily concentrated in the 1.0 to 1.5 range (Figure 5), indicating that lower severity levels are the most common. Frequency then gradually decreases as scores increase, with relatively few cases above 4.0. This suggests that mild severity is predominant in the dataset. We remove corrupted files (e.g., missing metadata) and non-English videos during preprocessing. The dataset is split into train/test sets, maintaining approximately a 3:1 ratio with stratified sampling based on task labels and downsampling of dominant categories to improve balance.
As shown in Table 8, the primary category task contains 1,966 training and 588 test samples. Training data exhibits natural class imbalance reflecting real-world distribution, ranging from 489 samples (“Relationship to Body”) to 66 samples (“Relationship to Masculinity”), while the test set maintains balanced representation (98 samples per category) for fair evaluation. The subcategory task operates on a filtered subset of 1,472 training and 462 test samples (Table 9). Training samples range from 200 (“Irrelevant”) to 28 (“Predebting Exercise”), while test data maintains balanced distribution for major categories (34 samples each) with reduced representation for rare categories (4-10 samples).
@@
Primary Category Train Test
Relationship to Body 489 98
Irrelevant 450 98
Relationship to Exercise 450 98
Relationship to Food 310 98
Supplement Abuse 201 98
Relationship to Masculinity 66 98
Total 1,966 588
Table 8: Distribution of samples across primary categories (Task 1).
Subcategory Train Test
Muscularity Self-objectification 170 34
Leanness Self-objectification 59 34
Muscle Dissatisfaction 72 34
Rigid Food Rules 144 34
Unsafe Food 59 34
Cheat Meals 60 34
Excessive Exercise 129 34
Predebting Exercise 28 4
Maladaptive Coping 82 34
Exercise-Induced Functional Impairment 43 6
Toxic Motivation 102 34
Anabolic Steroids 61 34
Legal APEDs 56 10
Hormone Therapy 77 34
Relationship to Masculinity 130 34
Irrelevant 200 34
Total 1,472 462
Table 9: Distribution of samples across subcategories (Task 2).
@@ D.1 Experiment Setup
Qwen2.5-VL 2025-06-05 7, 32, 72 Open
Table 10: Model version dates, parameter counts, and types (API vs. Open-source) for video-based VLMs.
@@ D.2 Prompt Templates
An example format of prompt used for zero-shot inference/supervised finetuning and few-shot inference is shown in Figures 6 and 7 and Figure 8 and 9:
@@
A11 6 occupational therapist specializing in functional rehabilitation and psychosocial support for eating-disorder patients in a neuropsychiatric hospital setting. Female
@@
Primary Secondary Definition Keywords
Relationship to Body Muscularity Self-Objectification Idealized imagery emphasizing muscular aesthetics as the primary source of value. shredded, swole, mensphysique, bodybuilding
Leanness Self-Objectification Self-evaluation against a lean, low-fat, highly-toned ideal promoted online. flat tummy, skinny men physique, small waist fitness
Muscle Dissatisfaction Expressing perceived insufficient muscularity despite fitness, fueling negative self-view. muscles never big enough, muscles not big enough, muscle dysmorphia, bdd men
Exercise-Induced Functional Impairment Prioritizing exercise over essential duties, harming daily functioning. gym over everything, gym or nothing, gym over friends, skip school for gym
Supplements Anabolic Steroids Normalization or endorsement of anabolic-androgenic steroid use with downplayed risks. tren, anabolic stack, anabolic gear, steroids
Relationship to Masculinity Links muscle-building and exercise performance to male identity, sexuality, and self-worth. gym masculinity, be a man gym, embrace masculinity, alpha male gym
Table 12: Primary and secondary categories of the Harm Taxonomy for Pro-Bigorexia Content on TikTok, with definitions and search keywords.
@@
7201534759347014958 Relationship to Masculinity #menshealth #menshealthtips #cynic hermes.the.cynic The reason that there is such a void of male role models for young men is because we literally do not have adult males in our population anymore. What do I mean by that? There’s no adult males. Like, there’s grown men everywhere, but they’re not actually physiologically adult human males…
7275865970306977067 Supplement Abuse this is totally healthy right? #recipes #creatine @days.of.j So I told a co-worker I was just taking my creatine in water, just water, because it was unflavored. And they told me you’re supposed to take it with like a meal or like sugar so then you can absorb it better…
7363842693211917611 Relationship to Food Liver King dinner! @liverking Pulled beef with hard shelled eggs Mmm It’s crazy. It actually tastes like the delicious pulled beef That’s in with the garlic oh and the onion and the lime juice booyah, yeah Sweet potatoes, the purple variety…
@@
7498781444408429867 Irrelevant Two in #shorts #gymshorts @johnbshop0 ugly, ugly, ugly aura. Bro, if your shorts, especially if you’re hitting leg day, are below your knees, you’re genuinely cooked, bro. These, I got them here on the TikTok shop. This is a size medium. I’m five foot 10, 165 pounds, and these things fit so good…
@@
7332272427986111790 Unsafe Food Liver King Dinner to start the week liverking Mmm, steak and potatoes, sweet potato fritters, Liver King Chef Lionel, that’s how he calls them. So good. We got…
@@
Table 14: Example entries for Subtype classification showing captions and TikTok metadata (Part 1 of 2). The actual videos can be found in the data attachment in the submission platform.
@@
7231695341077138730 Maladaptive Coping Gym Therapy #gymtherapy #gym #doordie #dobetter #military @shreddedvets Gym therapy is a term that was coined by Shredded Vets that describes the psychological effects of working out consistently…
7460046778524617991 Exercise-Induced Functional Impairment Bradman Best - I would skip school just to go to the gym. I was going to do whatever it took. I would skip school to guru the gym. lucky enough like when I was in year 9 and year 10 the Knights were giving me a day off…
@@
7313287497650326816 Anabolic Steroids #samsulek POV: Steroids gives you acne HAVE FUCKING I MAN LIKE PIMPLE SHIT COLLAB MIGHT PIMPLE UH ASIAN dude I think I have like a fucking I got like a cyst man you know like where it’s a…
@@
Table 15: Example entries for Subtype classification showing captions and TikTok metadata (Part 2 of 2). The actual videos can be found in the data attachment in the submission platform.

### v3

@@ A.1 Categories Definition
Throughout taxonomy development, clinical experts engage in extensive deliberation over edge cases that reveal the central challenge of drawing the line between harmful and legitimate fitness content—a distinction requiring specialized clinical expertise, real-world patient experience, and collaborative consensus-building. For instance, the Maladaptive Coping subcategory sparks debate about whether content promoting exercise as emotional regulation constitutes harm, with experts weighing clinical distinctions between healthy stress management and problematic avoidance behaviors. Similarly, Legal APEDs discussions center on sarcastic supplement content, where experts consider whether adult humor might inadvertently normalize excessive consumption among adolescent audiences. These deliberations require experts to balance contextual factors, including target audience, creator intent versus potential impact, and clinical thresholds between enthusiasm and pathology. Through structured consensus sessions examining ambiguous cases, annotators establish decision rules that balance clinical rigor with practical annotation consistency.
Primary Secondary Definition Keywords
Relationship to Body Muscularity Self‐Objectification Idealized imagery emphasizing muscular aesthetics as the primary source of value. shredded, swole, mensphysique, bodybuilding
Leanness Self‐Objectification Self‐evaluation against a lean, low‐fat, highly‐toned ideal promoted online. flat tummy, skinny men physique, small waist fitness
Muscle Dissatisfaction Expressing perceived insufficient muscularity despite fitness, fueling negative self‐view. muscles never big enough, muscles not big enough, muscle dysmorphia, bdd men
Exercise‐Induced Functional Impairment Prioritizing exercise over essential duties, harming daily functioning. gym over everything, gym or nothing, gym over friends, skip school for gym
Supplements Anabolic Steroids Normalization or endorsement of anabolic‐androgenic steroid use with downplayed risks. tren, anabolic stack, anabolic gear, steroids
Relationship to Masculinity Links muscle‐building and exercise performance to male identity, sexuality, and self‐worth. gym masculinity, be a man gym, embrace masculinity, alpha male gym
Table 7: Primary and secondary categories of the Harm Taxonomy for Pro‑Bigorexia Content on TikTok, with definitions and search keywords.
@@ B.1 Annotator Profiles
We enlist 16 subject‐matter experts, ranging from clinical psychologists and social workers to computational social scientists, for video annotation. Our annotators are research collaborators and co-authors who volunteered their clinical expertise with full knowledge of the project’s scope and scientific objectives, ensuring informed participation in this sensitive content annotation task. Table 13 details each annotator’s ID, batch assignment, area of expertise, and gender.
@@ B.3 Annotation Platform
The HTML interface (Figure 2) presents each TikTok video with its caption and hashtags (if any). The form beneath the video prompts the annotator to mark if the video is: (1) Showing One’s Body (yes/no); includes (2) Commercial Sponsorship (yes/no); (3) 2 boxes for Type of harm (1 mandatory, 1 optional), consisting of primary categories and subcategories, plus “Relevant but Not Listed,” “Irrelevant,” and “Unsure—Needs Consultation” options; and (4) Severity of Harm on a discrete scale ranging from 1 (Not Harmful) to 5 (Extremely Harmful). A text box allows annotators to write additional comments and edge‐case observations.
@@ B.4 Annotation Processing
After Round 2, we apply a five‐step pipeline to prepare the data for consensus review and final curation:
@@
    Data Filtering: remove videos that are unavailable (deleted) or non‐English, as indicated in annotators’ notes.
@@
    Reliability Assessment: compute Cohen’s $\kappa$ for both strong (exact primary–secondary match) and weak (primary‐only match) agreement to quantify inter‐rater reliability.
@@
    Consensus Review and Re‐annotation: filter out all items by agreement level (strong, weak, disagreement), sample a subset of weak and disagreement cases for each annotator pair, hold a batch‐specific meeting to discuss and jointly re‐annotate selected cases to reach consensus, then return the remaining weak and disagreement items to annotators for independent re‐annotation.
@@ B.5 Annotation Statistics
Inter-annotator agreement substantially improves from Round 2 to Final (Tables 3 and 4), with Cohen’s $\kappa$ values increasing from moderate agreement (0.43-0.69 strong $\kappa$, 0.59-0.83 weak $\kappa$) to good-to-excellent agreement (0.58-0.81 strong $\kappa$, 0.78-0.94 weak $\kappa$). The consistently high weak $\kappa$ values (>0.9 in most batches) in the Final round indicate that annotators achieved near-perfect agreement on the broader annotation categories.
@@
Table 8: Cohen’s $\kappa$ coefficients for inter-annotator agreement across annotation rounds. Strong $\kappa$ denotes exact agreement on the Subcategory label; Weak $\kappa$ denotes agreement on the Primary Category only. Reliability improved substantially in the Final round following consensus adjudication.
@@
(b) Subcategories
Table 9: Distribution of raw individual annotations (before consensus aggregation). Since each video received two independent annotations, these counts reflect the total number of labels assigned by annotators, whereas Table 1 reflects the final unique video counts after aggregation.
@@ Appendix C Data
Our dataset comprises TikTok videos annotated for hierarchical classification across two levels of granularity. The primary category classification categorizes content into six broad categories related to body image and health behaviors, while the subcategory classification provides fine-grained classification into 16 specific subcategories. The severity score scale defines the intensity of the harm emotion. The severity scores are most heavily concentrated in the 1.0 to 1.5 range (Figure 5), indicating that lower severity levels are the most common. The frequency then gradually decreases as scores increase, with relatively few cases exceeding 4.0. This suggests that mild severity is predominant in the dataset. Figure 7 shows that most content falls in the moderate-harm range, though Supplement Abuse and Relationship to Food score higher on average, likely because markers such as extreme dieting practices or steroid use are more readily identifiable to annotators.
We remove corrupted files (e.g., missing metadata) and non-English videos during preprocessing. The dataset is split into train/test sets, maintaining approximately a 3:1 ratio with stratified sampling based on task labels and downsampling of dominant categories to improve balance.
As shown in Table 1, the primary category task contains 1,966 training and 588 test samples. Training data exhibits natural class imbalance reflecting real-world distributions, ranging from 489 samples (“Relationship to Body”) to 66 samples (“Relationship to Masculinity”), while the test set maintains a strictly balanced representation (98 samples per category) for fair evaluation. The subcategory task operates on a filtered subset of 1,472 training and 462 test samples. Training samples range from 200 (“Irrelevant”) to 28 (“Predebting Exercise”), while test data maintains a balanced distribution for major categories (34 samples each) with reduced representation for rare categories (4–10 samples).
@@
Figure 6: Histogram of severity score differences between annotators. Positive values indicate higher severity by Annotator A. Colors denote agreement levels: green (exact), yellow ($\pm$1 point), red (>>2 points).
Figure 7: Box plot of severity scores across the primary types. Each box shows the distribution of annotated severity scores, with mean markers included. Dashed vertical lines indicate the harm thresholds, distinguishing low-harm content ($\leq$ 1.5) from high-harm content ($\geq$ 3.5)
@@ D.1 Experiment Setup
Qwen2.5‑VL 2025-06-05 7, 32, 72 Open
Table 10: Model version dates, parameter counts, and types (API vs. Open-source) for video‐based VLMs.
@@ D.2 Prompt Templates
An example format of prompt used for zero-shot inference/supervised finetuning and few-shot inference is shown in Figures 8 and 9 and Figure 10 and 11.
Appendix E Results
E.1 Hyperparameter Robustness Check
To validate our choice of using default temperature settings ($T=1.0$) for commercial API models, we re-evaluated Task 1 (Primary Category Classification) using a deterministic setting ($T=0.1$). Table 11 compares the performance across both settings.
The results demonstrate that performance remains highly stable, with F1 score variations confined within a narrow range ($\pm 0.01$). For instance, Claude-Sonnet-4 shows a slight improvement at $T=0.1$ (+0.008 F1), while GPT-4.1 shows a minor decrease (-0.004 F1). These marginal differences indicate that the specific temperature choice for high-performing commercial models does not significantly alter the comparative conclusions drawn in the main paper.
Model Temp Acc. Pm Rm F1m
Claude-Sonnet-4 1.0 0.829 0.832 0.829 0.827
Claude-Sonnet-4 0.1 0.837 0.841 0.837 0.835
GPT-4.1 1.0 0.796 0.808 0.792 0.792
GPT-4.1 0.1 0.787 0.809 0.787 0.788
Gemini-2.5-Flash 1.0 0.805 0.807 0.805 0.805
Gemini-2.5-Flash 0.1 0.810 0.813 0.810 0.809
Table 11: Robustness check for commercial API models on Task 1 (Primary Category) comparing default ($T=1.0$) vs. deterministic ($T=0.1$) temperature settings. $P_{m}$, $R_{m}$, and $F1_{m}$ denote macro precision, recall, and F1.
E.2 Per-Subtype Detection Analysis
Table 12 details the F1 scores for all 15 subcategories, revealing significant performance variance driven by content explicitness.
Explicitness Hierarchy
Subcategories characterized by distinct lexical keywords or unambiguous visual objects achieve the highest detection rates. Cheat Meals (Claude-4 ZS: 0.899) and Supplement Abuse subtypes (e.g., Anabolic Steroids, Hormone Therapy) consistently achieve F1 scores above 0.80, benefiting from clear indicators like specific drug names (e.g., “tren”, “TRT”). Conversely, performance drops sharply for ambiguous categories requiring subjective behavioral interpretation, such as Muscularity Self-objectification and Excessive Exercise (<0.60<0.60 F1), reflecting the difficulty in distinguishing harmful behaviors from normative training without deep context.
Model-Specific Strengths
Commercial models demonstrate superior reasoning capabilities on psychologically nuanced categories, with Claude-4 (Few-Shot) achieving the highest score on Maladaptive Coping (0.787). In contrast, open-source models like InternVL3-38B excel in categories with strong visual or explicit signals, leading performance in Unsafe Food and Anabolic Steroids (Few-Shot: 0.831), confirming their utility for detecting overtly harmful content.
Commercial Models Open-Source Models
Claude-4 Gemini-2.5 GPT-4.1 InternVL3-38B Qwen2.5VL-32B
Subcategory ZS FS ZS ZS FS ZS FS FT ZS FS FT
Muscularity Self-objectification 0.451 0.431 0.512 0.562 0.495 0.578 0.384 0.349 0.538 0.391
Leanness Self-objectification 0.676 0.635 0.645 0.706 0.636 0.590 0.500 0.533 0.625 0.561
Muscle Dissatisfaction 0.655 0.738 0.774 0.714 0.765 0.677 0.765 0.654 0.676 0.754 0.806
Rigid Food Rules 0.721 0.698 0.714 0.642 0.618 0.769 0.750 0.703 0.607 0.725 0.730
Unsafe Food 0.812 0.848 0.833 0.800 0.812 0.889 0.857 0.899 0.691 0.759 0.879
Cheat Meals 0.899 0.781 0.866 0.742 0.750 0.836 0.889 0.700 0.712 0.901 0.818
Anabolic Steroids 0.706 0.735 0.750 0.765 0.730 0.781 0.831 0.781 0.754 0.825 0.690
Legal APEDs 0.700 0.818 0.667 0.818 0.800 0.833 0.696 0.667 0.556 0.636 0.696
Hormone Therapy 0.769 0.794 0.712 0.788 0.824 0.783 0.794 0.778 0.776 0.794 0.800
Excessive Exercise 0.449 0.510 0.408 0.507 0.538 0.433 0.500 0.600 0.267 0.459 0.476
Maladaptive Coping 0.712 0.787 0.655 0.560 0.759 0.489 0.471 0.607 0.419 0.490 0.635
Exercise-induced Functional Impairment 0.571 0.533 0.500 0.500 0.308 0.533 0.429 0.333 0.333 0.353
Toxic Motivation 0.607 0.552 0.603 0.571 0.594 0.563 0.523 0.448 0.533 0.508 0.516
Relationship to Masculinity 0.779 0.740 0.621 0.806 0.712 0.618 0.691 0.691 0.588 0.759 0.649
Irrelevant 0.513 0.470 0.629 0.475 0.532 0.559 0.568 0.658 0.527 0.545 0.743
Table 12: Per-Class F1 Scores for Subcategory Classification by Model and Training Paradigm. ZS = Zero-Shot, FS = Few-Shot, FT = FineTuned. Highest scores per subcategory in green, second-highest in cyan. Dashes indicate missing data.
@@ E.2 Per-Subtype Detection Analysis
A11 6 occupational therapist specializing in functional rehabilitation and psychosocial support for eating‐disorder patients in a neuropsychiatric hospital setting. Female
@@
002 Relationship to Masculinity #menshealth #menshealthtips #cynic @<redacted> The reason that there is such a void of male role models for young men is because we literally do not have adult males in our population anymore. What do I mean by that? There’s no adult males. Like, there’s grown men everywhere, but they’re not actually physiologically adult human males…
003 Supplement Abuse this is totally healthy right? #recipes #creatine @<redacted> So I told a co-worker I was just taking my creatine in water, just water, because it was unflavored. And they told me you’re supposed to take it with like a meal or like sugar so then you can absorb it better…
004 Relationship to Food <redacted> dinner! @<redacted> Pulled beef with hard shelled eggs Mmm It’s crazy. It actually tastes like the delicious pulled beef That’s in with the garlic oh and the onion and the lime juice booyah, yeah Sweet potatoes, the purple variety…
@@
006 Irrelevant Two in #shorts #gymshorts @<redacted> ugly, ugly, ugly aura. Bro, if your shorts, especially if you’re hitting leg day, are below your knees, you’re genuinely cooked, bro. These, I got them here on the TikTok shop. This is a size medium. I’m five foot 10, 165 pounds, and these things fit so good…
@@
011 Unsafe Food Liver King Dinner to start the week <redacted> Mmm, steak and potatoes, sweet potato fritters, Liver King Chef Lionel, that’s how he calls them. So good. We got…
@@
015 Maladaptive Coping Gym Therapy #gymtherapy #gym #doordie #dobetter #military @<redacted> Gym therapy is a term that was coined by Shredded Vets that describes the psychological effects of working out consistently…
016 Exercise-Induced Functional Impairment @<redacted> - I would skip school just to go to the gym. I was going to do whatever it took. I would skip school to guru the gym. lucky enough like when I was in year 9 and year 10 the Knights were giving me a day off…
@@
018 Anabolic Steroids #<redacted> POV: Steroids gives you acne HAVE FUCKING I MAN LIKE PIMPLE SHIT COLLAB MIGHT PIMPLE UH ASIAN dude I think I have like a fucking I got like a cyst man you know like where it’s a…
@@
Table 15: Example entries for Subtype classification showing captions and TikTok metadata. The actual videos can be found in the data attachment in the submission platform.